what is the pattern rule 2 11 38 119
Answer by Edwin McCravy(20056)

an = 6n³ - 27n² + 48n - 25

a1 = 6(1)³-27(1)²+48(1)-25 = 6(1)-27(1)+48-25 = 6-27+48-25 = 2

a2 = 6(2)³-27(2)²+48(2)-25 = 6(8)-27(4)+96-25 = 48-108+96-25 = 11

a3 = 6(3)³-27(3)²+48(3)-25 = 6(27)-27(9)+144-25 = 162-243+144-25 = 38

a4 = 6(4)³-27(4)²+48(4)-25 = 6(64)-27(16)+192-25 = 384-432+192-25 = 119